OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, –4) with slope 4.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

point-slope formula y-y1=m(x-x1) where m=-4 x1=4 y1=1 and we have y-1=-4(x-4)

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = 4 (4,-4)..x1 = 4 and y1 = -4 now we just sub our stuff into the formula.. y - (-4) = 4(x - 4) y + 4 = 4(x - 4) <=== point slope form